Writers' Company produces 2 products presently
Fountain
Pen ball point
Unit 10000 20000
Area occupied (sq feet) 7000 8000
Variable cost - direct 32 25
Fixed cost - direct (per unit) 10 8

Rent – fixed cost per
Month(indirect – common for all products)
The Fixed indirect cost of Rent of Rs 3,00,000/- is to be allocated to both the products
The company was doing this on the "number of units" basis till now.
a. Please compute cost per unit of each product basis the above method.
1
Expert's answer
2020-12-07T07:32:53-0500

Writers' Company

                          Fountain’s cost per unit 

Total variable cost = "variable cost per unit \times units" = "Rs 32\times 10000 units" = "Rs 320,000"

Total fixed cost = "fixed cost per unit \times units = Rs 10 \times 10000 units = Rs 100,000"

Total Fountain’s rent = 

Total products’ units = 10000 + 20000 = 30000

Total Fountain’s rent = "\frac{10000}{30000} \times Rs 300,000 = Rs 100,000"

Total cost = total variable cost + total fixed cost + total rent

= Rs 320,000 + Rs 100,000 + Rs 100,000

= Rs 520,000

Cost per unit = "\frac{Total cost}{total units} = \frac{ Rs 520,000}{10000 units } =Rs 52"


Pen ball point’s cost per unit

Total variable cost = "variable cost per unit \\times \tunits = Rs 25 \\times 20000 units = Rs 500,000"

"Total fixed cost = fixed cost per unit \times units = Rs 8 \times 20000 units = Rs 160,000"

Total Pen ball point’s rent = 

"\frac{ Pen ball point’s units }{ total products’ units } \times the total fixed indirect cost of rent"

Total products’ units = 10000 + 20000 = 30000

Total Pen ball point’s rent = "\frac{ 20000}{ 30000} \times Rs 300,000 = Rs 200,000"

Total cost = total variable cost + total fixed cost + total rent

= Rs 500,000 + Rs 160,000 + Rs 200,000

= Rs 860,000

"Cost per unit = \frac{Total cost }{ total units } = \frac{ Rs 860,000}{ 20000 units } = Rs 43"